OpenStudy (anonymous):

The product of two positive numbers is 64. Find the numbers that have the least sum.

OpenStudy (anonymous):

x*y = 64 try having both x and y be the same number, so that their product is 64 and the sum is 16

OpenStudy (anonymous):

xy=64 x+y=s y=64/x x+(64/x)=s x+(64x^-1)=s 1-64x^-2=s 1-64x^-2=0 x=+-8

OpenStudy (anonymous):

-8 is not within are given domain or field therefore 8 is your answer.... so 8=x y=8

OpenStudy (anonymous):

from 1-64x^-2=s how did you jump to 1-64x^-2=0 would u tell me Outkast3r09

OpenStudy (anonymous):

from 1-64x^-2 is the derivative a maximum or minimum occurs when the derivative is equal to zero

OpenStudy (anonymous):

it should read 1-64x^-2=s' but i forgot the prime symbol